OpenStudy (anonymous):

Write the expression in standard form. 7/4-14i

OpenStudy (anonymous):

\[\frac{7}{4-14i}\] \[=\frac{7}{4-14i}\times \frac{4+14i}{4+14i}\] \[=\frac{7(4+14i)}{4^2+14^2}\] is a start